OpenStudy (anonymous):

simplify (25/16) to the -3/2 power

OpenStudy (ajprincess):

\[(\frac{25}{16})^\frac{-3}{2}\] \[=\frac{25^\frac{-3}{2}}{16^\frac{-3}{2}}\] \[=\frac{16^\frac{3}{2}}{25^\frac{3}{2}}\] \[=\frac{(4^2)^\frac{3}{2}}{(5^2)^\frac{3}{2}}\] Can u do nw?

OpenStudy (ajprincess):